# Data-driven Implementations of Various Generalizations of Balanced Truncation

### v2

@@ 3.2 RADI-based Non-intrusive Implementation of PR-BT
From Theorem 3.1, the RADI-based low-rank approximation of $P_{\text{PR}}$ can be obtained as follows. The projection matrix
@@
According to Theorem 3.1, the state-space realization of the ROM $\hat{G}_{\text{PR}}(s)=\hat{C}_{\text{PR}}(sI-\hat{A}_{\text{PR}})^{-1}\hat{B}_{\text{PR}}$ (where $\hat{A}_{\text{PR}}=S_{v}-Q_{v}^{-1}L_{v}^{T}L_{v}$, $\hat{B}_{\text{PR}}=Q_{v}^{-1}L_{v}^{T}$, and $\hat{C}_{\text{PR}}=C_{\text{PR}}V_{\text{PR}}$) solves the following projected Riccati equations with $\hat{P}_{\text{PR}}=Q_{v}^{-1}>0$:
@@
where $\hat{V}$ is as in (1) and $\hat{W}$ is an unknown projection matrix satisfying $\hat{W}^{*}\hat{V}=I$. According to Theorem 3.3, the ROM $\hat{G}_{\text{PR}}(s)$ in (46) interpolates $G_{\text{PR}}(s)$ at the interpolation points $(\sigma_{1},\cdots,\sigma_{v})$ since $V_{\text{PR}}=\hat{V}T_{v}$. An immediate consequence is that a ROM $\hat{G}_{\text{PR}}(s)$ interpolating $G_{\text{PR}}(s)$ be constructed directly from the ROM $\hat{G}(s)=C\hat{V}(sI-S_{v}+\zeta L_{v})^{-1}\zeta+D$ using the formula (46). We now use this observation to derive an expression for $T_{v}$.
@@
Next, the RADI-based approximation of $Q_{\text{PR}}$ can be obtained using Theorem 3.2 as follows. The projection matrix
@@
According to Theorem 3.2, the state-space realization of the ROM $\hat{G}_{\text{PR}}(s)=\hat{C}_{\text{PR}}(sI-\hat{A}_{\text{PR}})^{-1}\hat{B}_{\text{PR}}$ (where $\hat{A}_{\text{PR}}=S_{w}-L_{w}L_{w}^{T}P_{w}^{-1}$, $\hat{B}_{\text{PR}}=W_{\text{PR}}^{*}B_{\text{PR}}$, and $\hat{C}_{\text{PR}}=L_{w}^{T}P_{w}^{-1}$) satisfies the projected Riccati equations with $\hat{Q}_{\text{PR}}=P_{w}^{-1}>0$
@@
$\hat{A}^{*}\hat{Q}_{\text{PR}}+\hat{Q}_{\text{PR}}\hat{A}+(\hat{C}-\hat{B}^{*}\hat{Q}_{\text{PR}})^{*}R_{\text{PR}}^{-1}(\hat{C}-\hat{B}^{*}\hat{Q}_{\text{PR}}))=0.$ (54)
Theorem 3.4.
Let $G(s)$ be a positive-real transfer function. Define $\tilde{L}_{w}^{*}=R_{\text{PR}}^{-\frac{1}{2}}(L_{w}^{T}-\hat{B}_{\text{PR}}^{*})=\begin{bmatrix}l_{w,1}^{*}&\cdots&l_{w,q}^{*}\end{bmatrix}$ and $T_{w}=\text{blkdiag}(l_{w,1}^{*},\cdots,l_{w,q}^{*})$. Assume that:
@@
The proof is similar to the proof of Theorem 3.3 and hence omitted for brevity. ∎
@@
$(S_{w}-L_{w}\zeta-\hat{W}^{*}BR_{\text{PR}}^{-1}\zeta)^{*}T_{w}-T_{w}S_{w}^{*}+\zeta^{*}R_{\text{PR}}^{-\frac{1}{2}}L_{w}^{T}=0.$
@@
Similar to DD-ADI-BT [2], the RADI-based non-intrusive implementation of PR-BT can be directly obtained using the results of Theorems 3.3 and 3.4. The pseudo-code for this implementation, named “Data-driven PR-BT (DD-PR-BT)”, is provided in Algorithm 2.
RADI shifts for approximating $P_{\text{PR}}$: $(-\sigma_{1},\cdots,-\sigma_{v})$; RADI shifts for approximating $Q_{\text{PR}}$: $(-\mu_{1},\cdots,-\mu_{w})$; Data: $\big(H(\sigma_{1}),\cdots,H(\sigma_{v}),H(\mu_{1}),\cdots,H(\mu_{w}),G(\infty)\big)$ and $H^{\prime}(\sigma_{i})$ for $\sigma_{i}=\mu_{j}$; Reduced order: $r$.
Output: ROM: $(\hat{A},\hat{B},\hat{C},D)$
Algorithm 2 $D$-PR-BT
[…]
2: Set $D=G(\infty)$ and $R_{\text{PR}}=D+D^{T}$.
3: Set $S_{v}$, $L_{v}$, $S_{w}$, $L_{w}$ as in (6).
4: Compute $Q_{v}$ and $P_{w}$ by solving the Lyapunov equations (9) and (10).
[…]
6: Set $\hat{C}_{\text{PR}}=R_{\text{PR}}^{-\frac{1}{2}}C\hat{V}T_{v}$ and $\hat{B}_{\text{PR}}=T_{w}^{*}\hat{W}^{*}BR_{\text{PR}}^{-\frac{1}{2}}$.
7: Update $Q_{v}$ and $P_{w}$ by solving the Lyapunov equations (40) and (51).
8: Decompose $Q_{v}^{-1}$ and $P_{w}^{-1}$ as $Q_{v}^{-1}=\hat{L}_{p}\hat{L}_{p}^{*}$ and $P_{w}^{-1}=\hat{L}_{q}\hat{L}_{q}^{*}$.
[…]
10: Set $W_{r}=T_{w}\hat{L}_{q}\tilde{U}_{1}\tilde{\Sigma}_{r}^{-1/2}$ and $V_{r}=T_{v}\hat{L}_{p}\tilde{V}_{1}\tilde{\Sigma}_{r}^{-1/2}$.
11: Compute $\hat{A}$, $\hat{B}$, and $\hat{C}$ from (28).
3.3 RADI-based Non-intrusive Implementation of BR-BT
@@
By applying Theorem 3.1, the RADI-based approximation of $P_{\text{BR}}$ can be obtained as follows. The projection matrix
@@
According to Theorem 3.1, the ROM $(\hat{A}_{\text{BR}}=S_{v}-\hat{B}_{\text{BR}}^{\text{P}}L_{v},\hat{B}_{\text{BR}}^{\text{P}}=Q_{v}^{-1}L_{v}^{T},\hat{C}_{\text{BR}}^{\text{P}}=C_{\text{BR}}^{\text{P}}V_{\text{BR}})$ solves the following projected Riccati equation with $\hat{P}_{\text{BR}}=Q_{v}^{-1}>0$:
@@
$\hat{A}P_{\text{BR}}+\hat{P}_{\text{BR}}\hat{A}^{*}+\hat{B}\hat{B}^{*}+(\hat{P}_{\text{BR}}\hat{C}^{*}+\hat{B}D^{T})R_{p}^{-1}(\hat{P}_{\text{BR}}\hat{C}^{*}+\hat{B}D^{T})^{*}=0.$
Theorem 3.5.
Let $G(s)$ be a bounded-real transfer function. Define $\tilde{L}_{v}=D^{T}R_{p}^{-\frac{1}{2}}\hat{C}_{\text{BR}}^{\text{P}}+R_{b}^{\frac{1}{2}}L_{v}=\begin{bmatrix}l_{v,1}&\cdots&l_{v,p}\end{bmatrix}$ and $T_{v}=\text{blkdiag}(l_{v,1},\cdots,l_{v,p})$. Assume that:
@@
The proofs of 1 and 2 are similar to Theorem 3.3; hence, they are omitted for brevity.
@@
$(S_{v}-\zeta L_{v}+\zeta D^{T}R_{p}^{-1}C\hat{V})T_{v}-T_{v}S_{v}+\zeta R_{b}^{\frac{1}{2}}L_{v}=0.$
@@
As per Theorem 3.2, the RADI-based approximation of $Q_{\text{BR}}$ can be obtained as follows. The projection matrix
@@
According to Theorem 3.2, the ROM $(\hat{A}_{\text{BR}}=S_{w}-L_{w}\hat{C}_{\text{BR}}^{\text{Q}},\hat{B}_{\text{BR}}^{\text{Q}}=W_{\text{BR}}^{*}B_{\text{BR}}^{\text{Q}},\hat{C}_{\text{BR}}^{\text{Q}}=L_{w}^{T}P_{w}^{-1})$ solves the following projected Riccati equation with $\hat{Q}_{\text{BR}}=P_{w}^{-1}>0$:
$\hat{A}_{\text{BR}}^{*}\hat{Q}_{\text{BR}}+\hat{Q}_{\text{BR}}\hat{A}_{\text{BR}}+(\hat{C}_{\text{BR}}^{\text{Q}})^{*}\hat{C}_{\text{BR}}^{\text{Q}}+\hat{Q}_{\text{BR}}\hat{B}_{\text{BR}}^{\text{Q}}(\hat{B}_{\text{BR}}^{\text{Q}})^{*}Q_{\text{BR}}=0.$ (67)
@@
Theorem 3.6.
Let $G(s)$ be a bounded-real transfer function. Define $\tilde{L}_{w}^{*}=D^{T}R_{p}^{-\frac{1}{2}}\hat{B}_{\text{BR}}^{\text{Q}}+R_{c}^{\frac{1}{2}}L_{w}^{T}=\begin{bmatrix}l_{w,1}^{*}&\cdots&l_{w,q}^{*}\end{bmatrix}$ and $T_{w}=\text{blkdiag}(l_{w,1}^{*},\cdots,l_{w,q}^{*})$. Assume that:
@@
The proof is similar to that of Theorem 3.5 and hence omitted for brevity. ∎
@@
$\hat{B}_{\text{BR}}^{\text{Q}}$ $=W_{\text{BR}}^{*}B_{\text{BR}}^{\text{Q}}=\begin{bmatrix}G_{\text{BR}}^{\text{Q}}(\mu_{1})\\ \vdots\\ G_{\text{BR}}^{\text{Q}}(\mu_{w})\end{bmatrix}=\begin{bmatrix}\hat{G}_{\text{PR}}^{\text{Q}}(\mu_{1})\\ \vdots\\ \hat{G}_{\text{PR}}^{\text{Q}}(\mu_{w})\end{bmatrix}$
@@
Similar to DD-ADI-BT [2], the RADI-based non-intrusive implementation of BR-BT can be directly obtained using the results of Theorems 3.5 and 3.6. The pseudo-code for this implementation, named “Data-driven BR-BT (DD-BR-BT)”, is provided in Algorithm 3.
RADI shifts for approximating $P_{\text{BR}}$: $(-\sigma_{1},\cdots,-\sigma_{v})$; RADI shifts for approximating $Q_{\text{BR}}$: $(-\mu_{1},\cdots,-\mu_{w})$; Data: $\big(H(\sigma_{1}),\cdots,H(\sigma_{v}),H(\mu_{1}),\cdots,H(\mu_{w}),G(\infty)\big)$ and $H^{\prime}(\sigma_{i})$ for $\sigma_{i}=\mu_{j}$; Reduced order: $r$.
Output: ROM: $(\hat{A},\hat{B},\hat{C},D)$
Algorithm 3 $D$-BR-BT
[…]
2: Set $D=G(\infty)$, $R_{p}=I-DD^{T}$, $R_{b}=I+D^{T}R_{p}^{-1}D$, $R_{q}=I-D^{T}D$, and $R_{c}=I+DR_{q}^{-1}D^{T}$.
3: Set $S_{v}$, $L_{v}$, $S_{w}$, and $L_{w}$ as in (6).
4: Compute $Q_{v}$ and $P_{w}$ by solving the Lyapunov equations (9) and (10).
[…]
6: Set $\hat{C}_{\text{BR}}^{\text{P}}=R_{p}^{-\frac{1}{2}}C\hat{V}T_{v}$ and $\hat{B}_{\text{BR}}^{\text{Q}}=T_{w}^{*}\hat{W}^{*}BR_{q}^{-\frac{1}{2}}$.
7: Update $Q_{v}$ and $P_{w}$ by solving the Lyapunov equations (60) and (66).
8: Decompose $Q_{v}^{-1}$ and $P_{w}^{-1}$ as $Q_{v}^{-1}=\hat{L}_{p}\hat{L}_{p}^{*}$ and $P_{w}^{-1}=\hat{L}_{q}\hat{L}_{q}^{*}$.
[…]
10: Set $W_{r}=T_{w}\hat{L}_{q}\tilde{U}_{1}\tilde{\Sigma}_{r}^{-1/2}$ and $V_{r}=T_{v}\hat{L}_{p}\tilde{V}_{1}\tilde{\Sigma}_{r}^{-1/2}$.
11: Compute $\hat{A}$, $\hat{B}$, and $\hat{C}$ from (28).
3.4 ADI-based Non-intrusive Implementation of SW-BT

### v3

@@ 3.2 Non-intrusive ADI-based Approximation of PRBT
From Theorem 1, the RADI-based low-rank approximation of $P_{\text{PR}}$ can be obtained as follows. The projection matrix
@@
According to Theorem 1, the state-space realization of the ROM $\hat{G}_{\text{PR}}(s)=\hat{C}_{\text{PR}}(sI-\hat{A}_{\text{PR}})^{-1}\hat{B}_{\text{PR}}$ (where $\hat{A}_{\text{PR}}=S_{v}-Q_{v}^{-1}L_{v}^{T}L_{v}$, $\hat{B}_{\text{PR}}=Q_{v}^{-1}L_{v}^{T}$, and $\hat{C}_{\text{PR}}=C_{\text{PR}}V_{\text{PR}}$) solves the following projected Riccati equations with $\hat{P}_{\text{PR}}=Q_{v}^{-1}>0$:
@@
where $\hat{V}$ is as in (1) and $\hat{W}$ is an unknown projection matrix satisfying $\hat{W}^{*}\hat{V}=I$. According to Theorem 3, the ROM $\hat{G}_{\text{PR}}(s)$ in (46) interpolates $G_{\text{PR}}(s)$ at the interpolation points $(\sigma_{1},\cdots,\sigma_{v})$ since $V_{\text{PR}}=\hat{V}T_{v}$. An immediate consequence is that a ROM $\hat{G}_{\text{PR}}(s)$ interpolating $G_{\text{PR}}(s)$ can be constructed directly from the ROM $\hat{G}(s)=C\hat{V}(sI-S_{v}+\zeta L_{v})^{-1}\zeta+D$ using the formula (46). We now use this observation to derive an expression for $T_{v}$.
@@
Next, the RADI-based approximation of $Q_{\text{PR}}$ can be obtained using Theorem 2 as follows. The projection matrix
@@
According to Theorem 2, the state-space realization of the ROM $\hat{G}_{\text{PR}}(s)=\hat{C}_{\text{PR}}(sI-\hat{A}_{\text{PR}})^{-1}\hat{B}_{\text{PR}}$ (where $\hat{A}_{\text{PR}}=S_{w}-L_{w}L_{w}^{T}P_{w}^{-1}$, $\hat{B}_{\text{PR}}=W_{\text{PR}}^{*}B_{\text{PR}}$, and $\hat{C}_{\text{PR}}=L_{w}^{T}P_{w}^{-1}$) satisfies the projected Riccati equations with $\hat{Q}_{\text{PR}}=P_{w}^{-1}>0$
@@
$\hat{A}^{*}\hat{Q}_{\text{PR}}+\hat{Q}_{\text{PR}}\hat{A}+(\hat{C}-\hat{B}^{*}\hat{Q}_{\text{PR}})^{*}R_{\text{PR}}^{-1}(\hat{C}-\hat{B}^{*}\hat{Q}_{\text{PR}})=0.$ (54)
Theorem 4.
Let $G(s)$ be a positive-real transfer function. Define $\tilde{L}_{w}^{*}=R_{\text{PR}}^{-\frac{1}{2}}(L_{w}^{T}-\hat{B}_{\text{PR}}^{*})=\begin{bmatrix}l_{w,1}^{*}&\cdots&l_{w,w}^{*}\end{bmatrix}$ and $T_{w}=\text{blkdiag}(l_{w,1}^{*},\cdots,l_{w,w}^{*})$. Assume that:
@@
The proof is similar to the proof of Theorem 3 and hence omitted for brevity. ∎
@@
$(S_{w}-L_{w}\zeta-\hat{W}^{*}BR_{\text{PR}}^{-1}\zeta)^{*}T_{w}-T_{w}S_{w}^{*}+\zeta^{*}R_{\text{PR}}^{-\frac{1}{2}}L_{w}^{T}=0.$ (55)
@@
Similar to NI-ADI-BT [61], the ADI-based non-intrusive approximation of PRBT (NI-ADI-PRBT) can be directly obtained using the results of Theorems 3 and 4. Specifically, by computing Cholesky-like factorizations $Q_{v}^{-1}=\hat{L}_{p}\hat{L}_{p}^{*}$ and $P_{w}^{-1}=\hat{L}_{q}\hat{L}_{q}^{*}$, and defining the low-rank factors $\tilde{L}_{p}=\hat{V}T_{v}\hat{L}_{p}$ and $\tilde{L}_{q}=\hat{W}T_{w}\hat{L}_{q}$, the approximations admit the factorized representations $P_{\mathrm{PR}}\approx\tilde{L}_{p}\tilde{L}_{p}^{*}$ and $Q_{\mathrm{PR}}\approx\tilde{L}_{q}\tilde{L}_{q}^{*}$. Substituting the exact Gramian factors $L_{p},L_{q}$ in BSA with these approximations $\tilde{L}_{p},\tilde{L}_{q}$ yields NI-ADI-PRBT, which can be implemented using transfer function samples in the right half of the $s$-plane.
3.3 Non-intrusive ADI-based Approximations of BRBT
@@
By applying Theorem 1, the RADI-based approximation of $P_{\text{BR}}$ can be obtained as follows. The projection matrix
@@
According to Theorem 1, the ROM $(\hat{A}_{\text{BR}}=S_{v}-\hat{B}_{\text{BR}}^{\text{P}}L_{v},\hat{B}_{\text{BR}}^{\text{P}}=Q_{v}^{-1}L_{v}^{T},\hat{C}_{\text{BR}}^{\text{P}}=C_{\text{BR}}^{\text{P}}V_{\text{BR}})$ solves the following projected Riccati equation with $\hat{P}_{\text{BR}}=Q_{v}^{-1}>0$:
@@
$\hat{A}\hat{P}_{\text{BR}}+\hat{P}_{\text{BR}}\hat{A}^{*}+\hat{B}\hat{B}^{*}+(\hat{P}_{\text{BR}}\hat{C}^{*}+\hat{B}D^{T})R_{p}^{-1}(\hat{P}_{\text{BR}}\hat{C}^{*}+\hat{B}D^{T})^{*}=0.$
Theorem 5.
Let $G(s)$ be a bounded-real transfer function. Define $\tilde{L}_{v}=D^{T}R_{p}^{-\frac{1}{2}}\hat{C}_{\text{BR}}^{\text{P}}+R_{b}^{\frac{1}{2}}L_{v}=\begin{bmatrix}l_{v,1}&\cdots&l_{v,v}\end{bmatrix}$ and $T_{v}=\text{blkdiag}(l_{v,1},\cdots,l_{v,v})$. Assume that:
@@
The proofs of 1 and 2 are similar to Theorem 3; hence, they are omitted for brevity.
@@
$(S_{v}-\zeta L_{v}+\zeta D^{T}R_{p}^{-1}C\hat{V})T_{v}-T_{v}S_{v}+\zeta R_{b}^{\frac{1}{2}}L_{v}=0.$ (64)
@@
As per Theorem 2, the RADI-based approximation of $Q_{\text{BR}}$ can be obtained as follows. The projection matrix
@@
According to Theorem 2, the ROM $(\hat{A}_{\text{BR}}=S_{w}-L_{w}\hat{C}_{\text{BR}}^{\text{Q}},\hat{B}_{\text{BR}}^{\text{Q}}=W_{\text{BR}}^{*}B_{\text{BR}}^{\text{Q}},\hat{C}_{\text{BR}}^{\text{Q}}=L_{w}^{T}P_{w}^{-1})$ solves the following projected Riccati equation with $\hat{Q}_{\text{BR}}=P_{w}^{-1}>0$:
$\hat{A}_{\text{BR}}^{*}\hat{Q}_{\text{BR}}+\hat{Q}_{\text{BR}}\hat{A}_{\text{BR}}+(\hat{C}_{\text{BR}}^{\text{Q}})^{*}\hat{C}_{\text{BR}}^{\text{Q}}+\hat{Q}_{\text{BR}}\hat{B}_{\text{BR}}^{\text{Q}}(\hat{B}_{\text{BR}}^{\text{Q}})^{*}\hat{Q}_{\text{BR}}=0.$ (69)
@@
Theorem 6.
Let $G(s)$ be a bounded-real transfer function. Define $\tilde{L}_{w}^{*}=D^{T}R_{p}^{-\frac{1}{2}}\hat{B}_{\text{BR}}^{\text{Q}}+R_{c}^{\frac{1}{2}}L_{w}^{T}=\begin{bmatrix}l_{w,1}^{*}&\cdots&l_{w,w}^{*}\end{bmatrix}$ and $T_{w}=\text{blkdiag}(l_{w,1}^{*},\cdots,l_{w,w}^{*})$. Assume that:
@@
The proof is similar to that of Theorem 5 and hence omitted for brevity. ∎
@@
$\hat{B}_{\text{BR}}^{\text{Q}}$ $=W_{\text{BR}}^{*}B_{\text{BR}}^{\text{Q}}=\begin{bmatrix}G_{\text{BR}}^{\text{Q}}(\mu_{1})\\ \vdots\\ G_{\text{BR}}^{\text{Q}}(\mu_{w})\end{bmatrix}=\begin{bmatrix}\hat{G}_{\text{BR}}^{\text{Q}}(\mu_{1})\\ \vdots\\ \hat{G}_{\text{BR}}^{\text{Q}}(\mu_{w})\end{bmatrix}$
@@
Similar to NI-ADI-BT [61], the ADI-based non-intrusive approximation of BRBT (NI-ADI-BRBT) can be directly obtained using the results of Theorems 5 and 6. Specifically, by computing Cholesky-like factorizations $Q_{v}^{-1}=\hat{L}_{p}\hat{L}_{p}^{*}$ and $P_{w}^{-1}=\hat{L}_{q}\hat{L}_{q}^{*}$, and defining the low-rank factors $\tilde{L}_{p}=\hat{V}T_{v}\hat{L}_{p}$ and $\tilde{L}_{q}=\hat{W}T_{w}\hat{L}_{q}$, the approximations admit the factorized representations $P_{\mathrm{BR}}\approx\tilde{L}_{p}\tilde{L}_{p}^{*}$ and $Q_{\mathrm{BR}}\approx\tilde{L}_{q}\tilde{L}_{q}^{*}$. Substituting the exact Gramian factors $L_{p},L_{q}$ in BSA with these approximations $\tilde{L}_{p},\tilde{L}_{q}$ yields NI-ADI-BRBT, which can be implemented using transfer function samples in the right half of the $s$-plane.
3.4 Non-intrusive ADI-based Approximation of SWBT